# EMT: A Visual Multi-Task Benchmark Dataset for Autonomous Driving in the Arab Gulf Region

### v1

@@ Abstract
This paper introduces the Emirates Multi-Task (EMT) dataset - the first publicly available dataset for autonomous driving collected in the Arab Gulf region. The EMT dataset captures the unique road topology, high traffic congestion, and distinctive characteristics of the Gulf region, including variations in pedestrian clothing and weather conditions. It contains over 30,000 frames from a dash-camera perspective, along with 570,000 annotated bounding boxes, covering approximately 150 kilometers of driving routes. The EMT dataset supports three primary tasks: tracking, trajectory forecasting and intention prediction. Each benchmark dataset is complemented with corresponding evaluations: (1) multi-agent tracking experiments, focusing on multi-class scenarios and occlusion handling; (2) trajectory forecasting evaluation using deep sequential and interaction-aware models; and (3) intention benchmark experiments conducted for predicting agents’ intentions from observed trajectories. The dataset is publicly available at avlab.io/emt-dataset, and pre-processing scripts along with evaluation models can be accessed at github.com/AV-Lab/emt-dataset.
@@ I Introduction
As autonomous driving technology advances, the ability of data-driven models to generalize across diverse road environments and conditions is essential for safe operation, but remains a significant challenge. To achieve robust generalization, it is critical to train models on datasets that capture a wide range of traffic scenes and characteristics. Current autonomous driving datasets provide extensive coverage of regions like the USA [1, 2, 3, 4, 5], Europe [6, 7], and parts of Asia, including China and Singapore [8, 1]. However, the Arab Gulf region, with its unique driving conditions, remains underrepresented. To address this gap, we introduce the Emirates Multi-Task (EMT) dataset, collected in the Unighted Araba Emirates (UAE) to capture the region’s distinct traffic conditions. This region offers diverse driving challenges due to its range of road layouts, including expansive highways, urban areas, and complex city junctions. Additionally, driving behavior in the UAE reflects a blend of modern regulations and traditional practices. Our dataset was gathered using dash cameras mounted on two vehicles driving across the UAE’s largest cities and on intercity routes, covering over 150 kilometers. The annotated dataset supports multiple benchmarks, including tracking, trajectory prediction, and intention prediction, aimed at advancing models robustness in complex driving environments.
@@
The tracking benchmark dataset is designed to evaluate the ability of algorithms to accurately identify and maintain consistent object tracking over time in a complex driving environment. Similar to current state-of-the-art (SOTA) tracking benchmarks [9, 1, 10], it focuses on the motion of vehicles, pedestrians, cyclists, and motorbikes, captured from a frontal camera perspective. The benchmark is designed to test tracking models under varying levels of traffic congestion and frequent lane changes. The dataset contains 8,806 unique tracking IDs, including 8,076 vehicles, 568 pedestrians, 158 motorbikes and 14 cyclists, and with a mean tracking duration of 6.5 seconds.
@@
Accurate trajectory prediction is a cornerstone of safe autonomous driving, enabling vehicles to navigate complex interactions and adjust their planning accordingly. The designed trajectory prediction benchmark dataset challenges forecasting models to predict trajectories in heterogeneous traffic and to generalize effectively across scenarios involving interacting agents, forecasts in large intersections and roundabouts, and multi-agent dynamics. The dataset contains 4,821 unique agents, including 4,347 vehicles and 386 pedestrians. Details on the size of the dataset, splits into past trajectories and prediction horizons, as well as the shifting window approach, are provided in Section III.
@@
The intention prediction benchmark is designed to evaluate the ability of autonomous systems to anticipate the future actions of traffic participants. This benchmark enables systems to infer the likely movements of nearby agents based on their current trajectories, positions, and the surrounding context. The dataset includes detailed annotations for vehicle maneuver intentions and pedestrian behaviors. For vehicles, the labeled maneuvers comprise high-level intended actions that define their future trajectories, including: turn left/right, keep lane, merge left/right, brake, stop and reverse. For pedestrians, the dataset provides labels for four intention classes: waiting to cross, crossing, walking (on pavements or sidewalks), and stopping (e.g., waiting at bus stops). These annotations aim to provide a comprehensive understanding of pedestrian and vehicle behaviors in diverse scenarios. In total the dataset contains 4,921 sequences of agents trajectories with intentions.
Each task-specific dataset is complemented by evaluated models and conducted experiments. For multi-agent tracking, we evaluate Kalman filter-based SOTA trackers within a pipeline that includes three configurations: off-shelf detector, a fine-tuned detector, and ground-truth detections. This setup allows us to systematically assess the trackers’ abilities to handle occlusions and detector errors. For trajectory prediction, we selected three deep-learning architectures, focusing on their capacity to capture temporal dependencies and interaction dynamics in high-density scenarios. Lastly, for the intention prediction task, we conducted evaluation to assess LSTM-based models’ ability to predict future intentions based on past trajectories.
@@ II-A Autonomous Driving Datasets
Among intention prediction datasets, the NGSIM111 http://ops.fhwa.dot.gov/trafficanalysistools/ngsim. htm and HighD [12] datasets focus on highway scenarios. NGSIM provides intention labels for lane-changing, lane-keeping, merging, and yielding, while HighD includes labels for free-driving, vehicle following, critical maneuvers, and lane-changing. HighD offers high-resolution highway data with trajectories spanning approximately 13 seconds per vehicle. NuScenes [1] and Argoverse [3] datasets cover urban driving, providing multimodal sensor data. The NuScenes dataset includes intentions such as moving, stopped, and parked. Argoverse contains maneuvers for lane changes, lane keeping, and left/right turns. The INTERACTION dataset [13] focuses on challenging environments like roundabouts, intersections, and highways, providing intention labels for lane changes, merging, and yielding, which are essential for interaction-aware trajectory prediction. BLVD dataset [14] offers diverse behavior and interaction annotations for urban driving, capturing a wide range of maneuvers such as lane changing, lane keeping, accelerating, decelerating, stopping, and left/right turns. This dataset supports both tracking and maneuver intention tasks through multi-camera video data. Lastly, the InD dataset [15] includes intention labels for actions such as merging, yielding, and left/right turns, with annotations similar to BLVD for vehicles, pedestrians, and riders.
The designed EMT dataset complements existing datasets by introducing data collected from a previously underrepresented geographical location. By including diverse driving scenarios typical for the Arab Gulf-region traffic, the EMT dataset provides a unique context for improving models’ predictive accuracy in culturally and regionally distinct driving behaviors.
@@ III Dataset
TABLE I: Object classes description and statistics.
@@ III-B Data Annotations
Every significant object and actor within each frame is annotated, including vehicles, pedestrians, small motorized vehicles, motorbikes, and cyclists. For objects classification and description we relied on convention presented in [63]. Small motorized vehicles refer to any motorized vehicle smaller than a car, such as mobility scooters or quad bikes. Medium vehicles include those larger than a car, such as vans, while large vehicles refer to lorries, typically characterized by having six or more wheels. Emergency vehicles include ambulances, police cars, and fire engines equipped with red and blue flashing lights, excluding vehicles with yellow flashing lights, such as road maintenance vehicles. Each object is assigned an intention label based on its observed behavior. Table I provides for each class description, number of bounding boxes and number of agents throughout the whole dataset.
@@
The most common road topologies for the region are illustrated in Figure 3. One observed scenario involves bridges appearing above the ego vehicle in the images. Vehicles on these bridges are ignored as they do not influence the ego vehicle’s planning. Figure 3(a) depicts a typical large city junction common in the region. A key feature of such junctions, compared to those in other areas, is the inclusion of a free right turn to alleviate traffic congestion. Cases (3) and (4) highlight this scenario: vehicles can make a free right turn from the traffic lights if no pedestrians are present and the turn does not interfere with other traffic. Otherwise, they must wait to complete the turn and merge safely onto the intended road. For clarity, the figure also includes other typical scenarios: (1) a car waiting at the junction for the green light; (2) a bus crossing the junction from left to right while staying in the same lane; (3) a car performing a free right turn to merge into a lane; (4) a car slowing down to allow a pedestrian to cross, tagged with the action ”braking”; and (5) a pedestrian standing on the pavement, waiting for the car to stop before crossing safely. Figure 3(b) schematically depicts a roundabout with eight access points, showing vehicles in various stages: approaching, circulating inside, and exiting. When the ego vehicle intends to enter the roundabout, it must monitor agents in the outer or both lanes, depending on its intended exit, to avoid collisions. Each vehicle within the roundabout is labeled with an intention, such as ”entering,” ”waiting to enter,” ”moving within the roundabout,” or ”exiting,” to provide a clear understanding of its behavior. Lastly, Figure 3(c) represents a typical highway layout for the region. It illustrates cases where a vehicle exits the highway and gradually separates from the main road. When a vehicle begins exiting by curving right and changing lanes, its intention is labeled as ”merge-right.” Once the vehicle has exited, the label remains ”keep-lane” until the vehicle disappears from the ego vehicle’s field of vision. All scenarios are supplemented with corresponding frames from the dataset.
@@
TABLE II: Description of Collected Videos
Videos (Train/Test) Number of Bounding Boxes Mean Density of Agents per frame Resolution Location Lightning Conditions Weather Conditions Road Elements
video_115533 (train) 38,841 21.6 2560x1440 Abu Dhabi day clear driving in city, highways
video_115833 (train) 47,433 26.3 2560x1440
video_122233 (test) 27,194 15.1 2560x1440 leaving city, highways
video_141733 (train) 23,135 12.7 896x672 driving in city, junctions
video_161205 (train) 35,771 19.6 896x672
video_151901 (train) 29,107 22.2 3840x2160 evening clear driving in city, at parking
video_143739 (train) 46,894 26 3840x2160 straight road
video_142911 (train) 25,573 14.2 3840x2160 heavy rain straight road
video_155425 (test) 18,485 13.9 3840x2160 rainy driving in city, at roundabout, merging main road
video_160325 (test) 20,581 11.4 3840x2160 driving in city, straight road
video_054604 (train) 45,576 25.3 2560x1440 Dubai day clear driving in city, highways
video_054907 (train) 52,981 29.4 2560x1440 leaving city, highways
video_125233 (test) 31,882 17.7 2560x1440 urban area, highways
video_131333 (train) 21,873 13.5 2560x1440 urban area, highways, bridges
video_204347 (test) 12,283 6.8 2560x1440 evening clear driving in city, bridges
video_204647 (train) 9,752 5.4 2560x1440 highways, bridges
video_174340 (train) 24,249 13.4 2560x1440 driving in city, highways
video_214547 (train) 35,404 19.6 2560x1440 driving in city, highways, bridge
video_210906 (train) 15,963 8.86 2560x1440 driving in city
video_220047 (train) 13,397 7.4 2560x1440
TABLE III: Trajectory Prediction Settings: The notation 30/60 indicates the duration of the past trajectory and prediction horizon, measured in frames. This setup corresponds to 3 seconds of observed data followed by 6 seconds of predicted trajectory. Each row corresponds to different sliding window setting and shows the size of train and test datasets.
@@ III-C Annotations Format
"id": 12,
"class": "Large_vehicle",
"frames": [109, 110, 111, 112, 113],
"bbox": [
[1934.9579, 1036.7965, 1956.0014, 1054.1146],
[1921.8802, 1038.1977, 1942.9237, 1055.5157],
[1904.132, 1041.4671, 1925.1755, 1058.7852],
[1887.3178, 1041.4671, 1908.3613, 1058.7852],
[1872.839, 1040.5329, 1893.8825, 1057.851]
"intention": ["turn-right", "turn-right", "turn-right", "braking", "braking"]
To enhance usability, we supplement the annotations with a parsing script capable of generating custom settings by varying the size of past trajectories, prediction horizon, and overlaps. The script iterates through all objects and segments their trajectories into samples of past and future trajectories using a sliding window approach. For example, if the sliding window is set to 3 frames, the first sample will be generated starting from frame_1, and the next from frame_4. The default format is based on past observations and is structured as: $[past\_trajectory]\ [future\_trajectory]$. Table III presents various settings along with the sizes of the training and test datasets. Additionally, we provide a scene-centered generation option, with further details available in Section IV-B.
III-D EMT Statistics
The dataset contains 20 videos, encompassing both day and night conditions, as well as a variety of weather scenarios, including rainy and clear weather. Table II provides details for each video, including the city, time of driving, weather conditions, and road topologies. In total, the dataset contains approximately 570,000 bounding boxes. The Table II also reports the mean density of agents per frame for each video, ranging from a minimum of around 6 agents per frame to a maximum of 30 agents, highlighting the density and diversity of the dataset.
III-E Annotation Tool
The annotations were generated using the Basic.ai tool - annotation platform equipped with tracking algorithms, propagating bounding boxes across frames while maintaining consistent unique identifiers. Its robust tracking functionality was a key reason for its selection, significantly streamlining the annotation process. On average, each video required approximately 70 hours of annotation, completed by a professional human annotator.

### v3

@@ Abstract
This paper introduces the Emirates Multi-Task (EMT) dataset, designed to support multi-task benchmarking within a unified framework. It comprises over 30,000 frames from a dash-camera perspective and 570,000 annotated bounding boxes, covering approximately 150 kilometers of driving routes that reflect the distinctive road topology, congestion patterns, and driving behavior of Gulf region traffic. The dataset supports three primary tasks: tracking, trajectory forecasting, and intention prediction. Each benchmark is accompanied by corresponding evaluations: (1) multi-agent tracking experiments addressing multi-class scenarios and occlusion handling; (2) trajectory forecasting evaluation using deep sequential and interaction-aware models; and (3) intention prediction experiments based on observed trajectories. The dataset is publicly available at avlab.io/emt-dataset, with pre-processing scripts and evaluation models at github.com/AV-Lab/emt-dataset.
@@ I Introduction
TABLE I: Comparison with existing camera-based trajectory–intention datasets highlighting that EMT offers the most extensive set of annotated intentions for both vehicles and pedestrians, along with the most diverse range of traffic scenarios.
Dataset Region Tasks Traffic Participants Intention Labels Traffic Scenarios
HighD [1] Germany - vehicles free-driving, vehicle following, critical maneuver, lane-change highway
JAAD [2] USA Canada Eastern Europe Intention Prediction Behavior Detection pedestrians bicyclists motorcyclists other drivers crossing, stopped, moving-fast, moving-slow, speed-up, slow-down intersections urban
PIE [3] Canada Trajectory Prediction Intention Prediction Trajectory Prediction conditioned on Intention pedestrians cross, not-cross urban
STIP [4] USA Intention Prediction pedestrians cross, not-cross urban
EMT (Ours) UAE Detection Tracking Trajectory Prediction Intention Prediction vehicles pedestrians motorbikes cyclists turn-right, turn-left, merge-right, merge-left, braking, stopped, lane-keeping, reversing, walking, crossing, waiting-to-cross intersections roundabouts urban highway
As autonomous driving systems mature, the research community increasingly focuses on the robustness of autonomous vehicle behavior under failures or inaccuracies in different components of the autonomy stack [5, 6]. Two critical modules influencing both motion planning and overall system safety are perception and prediction, with the latter inherently dependent on the output of the former. Beyond the interdependence between modules, tasks within each module are also tightly coupled. In perception block, detection and tracking are interlinked, i.e., errors in detection propagate to tracking [7, 8, 9]. In the prediction module, trajectory forecasting and behavioral inference (e.g., pedestrian crossing intention or vehicle maneuvers) are strongly correlated with the tracking perfomance. These interdependencies underscore the need for datasets that support not only task-specific evaluation using ground truth inputs, but also multi-task evaluation where performance dependencies between tasks are preserved.
In addition to core tasks, recent datasets increasingly incorporate semantic attributes to bridge the gap between raw perception and high-level scene understanding. One such attribute is the intention or maneuver of traffic participants. Human drivers naturally interpret these cues, e.g., a turn signal or lateral drift suggests a lane change for a vehicle. These anticipations, formalized as intentions, shape trajectories and inform collision avoidance. Yet, intention-labeled datasets [1, 2, 3, 4, 10, 11] remain limited, including joint evaluation of intention and trajectory. Capturing this coupling can facilitate more anticipatory and improve trajectory forecasting accuracy.
Lastly, the ability of data-driven models to generalize across diverse road environments and conditions is essential for safe autonomous operation, yet remains a persistent challenge. Robust generalization requires training on datasets that span a wide range of traffic scenarios. However, the datasets remain scarce, since large-scale annotations are costly and require immense human labor, even with semi-automated annotation pipelines. Additionally, the existing datasets provide coverage of regions such as the USA [8, 12, 13, 14, 15], Europe [16, 17], and parts of Asia including China and Singapore [18, 8]. The distribution of collected remains geographically uneven, leaving a gap to collect the data from unrepresented regions.
To contribute toward addressing the discussed gaps, we introduce EMT (Emirates Multitask Dataset) — a dataset constructed from frontal camera footage and designed to support multi-task evaluation. EMT enables consistent benchmarking of object detection, tracking, trajectory forecasting, and intention prediction on temporally aligned data. This structure allows researchers to evaluate each task independently and to examine cross-task dependencies, such as how perception-level errors propagate and affect downstream modules. The dataset is collected in the United Arab Emirates (UAE), a region characterized by a wide range of driving scenarios spanning expansive highways, urban centers, and complex intersections. Specifically, EMT comprises the following task-specific datasets.
@@
The tracking benchmark dataset is designed to evaluate the ability of algorithms to accurately identify and maintain consistent object tracking over time in a complex driving environment for vehicles, pedestrians, cyclists, and motorbikes. The benchmark is designed to test tracking models under varying levels of traffic congestion and frequent lane changes. The dataset contains 8,806 unique tracking IDs, including 8,076 vehicles, 568 pedestrians, 158 motorbikes and 14 cyclists, and with a mean tracking duration of 6.5 seconds.
@@
The trajectory prediction benchmark dataset challenges forecasting models to predict trajectories in heterogeneous traffic and to generalize effectively across scenarios involving interacting agents, forecasts in large intersections and roundabouts, and multi-agent dynamics. The dataset contains 4,821 unique agents, including 4,347 vehicles and 386 pedestrians. Details on the size of the dataset, splits into past trajectories and prediction horizons, as well as the shifting window approach, are provided in Section III.
@@
The intention prediction benchmark dataset is designed to test models ability to infer the likely actions of surrounding agents based on their past trajectories. For vehicles, the dataset contains maneuver labels: turn left/right, keep lane, merge left/right, brake, stop and reverse and for pedestrians: waiting to cross, crossing, walking (on pavements or sidewalks), and stopping (e.g., waiting at bus stops). These annotations aim to provide a comprehensive understanding of pedestrian and vehicle behaviors. In total the dataset contains 4,921 sequences of agents trajectories with intentions.
Each task-specific dataset is accompanied by baseline models. For cross-task analysis, we include experiments demonstrating the effect of detection quality on tracking performance. Our key contributions are as follows: 1) a diverse visual dataset that supports multi-task benchmarking, enabling both per-task evaluation and analysis of tasks interdependencies critical for real-world deployment; 2) a benchmark for evaluating the generalization of models trained on data from other regions, offering the resource for assessing performance in underrepresented Gulf driving environments.
@@ II-A Autonomous Driving Datasets
In the context of multi-task learning, BDD100K [19] is a large-scale dataset comprising approximately 100K driving videos, designed to support multi-task training across 10 vision tasks relevant to autonomous driving, including object detection, tracking, semantic segmentation, lane marking, and drivable area detection. Among trajectory prediction datasets, nuScenes [8], Argoverse [13, 14], and WAYMO [9] motion forecasting datasets are commonly used for state-of-the-art model evaluation. Focusing on interactions between traffic participants, the INTERACTION dataset [20] captures complex environments such as roundabouts, intersections, and highways, offering scenarios critical for interaction-aware trajectory prediction.
Related to intention prediction datasets, Table I includes a comparison between the proposed EMT and existing camera-based intention datasets [1, 2, 3, 4]. In addition, among multi-sensor datasets, the BLVD dataset [10] was collected in China and designed primarily for tasks such as interactive event recognition and intention prediction. Events in BLVD combine landmark estimation with fine-grained intention prediction. The dataset covers vehicles, pedestrians, and riders (cyclists and motorcyclists), and provides detailed intention labels: 13 types for vehicles, 8 for pedestrians, and 7 for riders, enabled by the sensor fusion setup that allows for accurate velocity estimation. Similarly, LOKI dataset [11] was collected in Japan using a multi-sensor rig. The dataset targets long-term and short-term intention prediction in dense urban scenes. It provides both frame-wise labels (e.g., stop, lane change, cut-in, walk, wait) and long-term goal annotations vehicles and pedestrians.
The designed EMT dataset complements existing camera-based intention datasets by offering a more fine-grained set of intention labels and greater diversity in traffic scenarios. Among multi-sensor intention datasets, EMT also fills a geographic gap by contributing data collected from a previously underrepresented region. By capturing diverse driving conditions typical of Arab Gulf-region traffic, EMT introduces a unique context that supports the development of models with improved predictive accuracy under culturally and regionally distinct driving behaviors.
@@ III Dataset
TABLE II: Object classes description and statistics.
@@ III-B Data Annotations
Every significant object and actor within each frame is annotated, including vehicles, pedestrians, small motorized vehicles, motorbikes, and cyclists. For objects classification and description we relied on convention presented in [68]. Small motorized vehicles refer to any motorized vehicle smaller than a car, such as mobility scooters or quad bikes. Medium vehicles include those larger than a car, such as vans, while large vehicles refer to lorries, typically characterized by having six or more wheels. Emergency vehicles include ambulances, police cars, and fire engines equipped with red and blue flashing lights, excluding vehicles with yellow flashing lights, such as road maintenance vehicles. Each object is assigned an intention label based on its observed behavior. Table II provides for each class description, number of bounding boxes and number of agents throughout the whole dataset.
@@
TABLE III: Description of Collected Videos Train and Test Splits
Videos Number of Bounding Boxes Number of Frames Resolutions Lightning and Weather Conditions,           Road Elements
Train Set video_115533, video_115833, video_141733, video_161205, video_151901, video_143739, video_142911, video_054604, video_054907, video_131333, video_204647, video_214547, video_210906, video_220047, video_174340 471,649 26,360 2560x1440 896x672 3840x2160 day, evening clear weather, heavy rain driving in city, highways, junctions, at parking, at roundabouts, merging main road, bridges, between cities
Test Set video_122233, video_204347, video_125233, video_155425, video_160325 98,049 8,530 2560x1440 3840x2160 day, evening clear weather, rainy driving in city, highways, at roundabouts, merging main road, bridges, between cities
TABLE IV: Trajectory Prediction Settings: The notation 30/60 indicates the duration of the past trajectory and prediction horizon, measured in frames. This setup corresponds to 3 seconds of observed data followed by 6 seconds of predicted trajectory. Each row corresponds to different sliding window setting and shows the size of train and test datasets.
@@
The most common road topologies for the region are illustrated in Figure 3. Figure 3(a) depicts a typical large city junction common in the region. A key feature of such junctions, compared to those in other areas, is the inclusion of a free right turn to alleviate traffic congestion. Cases (3) and (4) highlight this scenario: vehicles can make a free right turn from the traffic lights if no pedestrians are present and the turn does not interfere with other traffic. Otherwise, they must wait to complete the turn and merge safely onto the intended road. For clarity, the figure also includes other typical scenarios: (1) a car waiting at the junction for the green light; (2) a bus crossing the junction from left to right while staying in the same lane; (3) a car performing a free right turn to merge into a lane; (4) a car slowing down to allow a pedestrian to cross, tagged with the action ”braking”; and (5) a pedestrian standing on the pavement, waiting for the car to stop before crossing safely. Figure 3(b) schematically depicts a roundabout with eight access points, showing vehicles in various stages: approaching, circulating inside, and exiting. When the ego vehicle intends to enter the roundabout, it must monitor agents in the outer or both lanes, depending on its intended exit, to avoid collisions. Each vehicle within the roundabout is labeled with an intention, such as ”stopped” if it is waiting to enter, ”keep-lane” in case of moving within the roundabout, and ”merge-right” when exiting the roundabout. Lastly, Figure 3(c) represents a typical highway layout for the region. It illustrates cases where a vehicle exits the highway and gradually separates from the main road. When a vehicle begins exiting by curving right and changing lanes, its intention is labeled as ”merge-right.” Once the vehicle has exited, the label remains ”keep-lane” until the vehicle disappears from the ego vehicle’s field of vision. All scenarios are supplemented with corresponding frames from the dataset.
@@ III-C Annotations Format
To enhance usability, we supplement the annotations with a parsing script capable of generating custom settings by varying the size of past trajectories, prediction horizon, and overlaps. The script iterates through all objects and segments their trajectories into samples of past and future trajectories using a sliding window approach. For example, if the sliding window is set to 3 frames, the first sample will be generated starting from frame_1, and the next from frame_4. The default format is based on past observations and is structured as: $[past\_trajectory]\ [future\_trajectory]$. Table IV presents various settings along with the sizes of the training and test datasets. Additionally, we provide a scene-centered generation option, with further details available in Section IV-B.
While all tasks are derived from the same video sequences and temporally aligned scenes, task-specific filtering is applied during preprocessing. For instance, parked vehicles are retained in the tracking benchmark to ensure that models learn to distinguish static objects and maintain consistent identities. In contrast, such objects are excluded from trajectory forecasting, which focuses on modeling agent dynamics. These differences necessitate task-specific object enumeration, resulting in non-shared object IDs across tasks. Nonetheless, the datasets share the same image streams and maintain consistent training and testing splits, ensuring contextual coherence for cross-task dependencies.
III-D EMT Statistics and Annotations
The dataset contains 20 videos, encompassing both day and night conditions, as well as a variety of weather scenarios, including rainy and clear weather. Table III provides details for each video, including the city, time of driving, weather conditions, and road topologies. In total, the dataset contains approximately 570,000 bounding boxes. The Table III also reports the mean density of agents per frame for each video, ranging from a minimum of around 6 agents per frame to a maximum of 30 agents, highlighting the density and diversity of the dataset.
The annotations were created using the Basic.ai tool — annotation platform equipped with tracking algorithms, propagating bounding boxes across frames while maintaining consistent unique identifiers. Its robust tracking functionality was a key reason for its selection, significantly streamlining the annotation process. On average, each video required approximately 70 hours of annotation, completed by professional human annotators.